Squids and octopuses propel themselves by expelling water. They do this by keeping water in a cavity and then suddenly contracting the cavity to force out the water through an opening. A 7.50 kg squid (including the water in the cavity) at rest suddenly sees a dangerous predator.

A. If the squid has 1.55 kg of water in its cavity, at what speed must it expel this water to suddenly achieve a speed of 2.60 m/s to escape the predator? Neglect any drag effects of the surrounding water.

B. How much kinetic energy does the squid create by this maneuver?
Answer #1

Solve it using the conservation of momentum principle.
(a) The total momentum is equal to zero since the squid was at rest in the begining. Now as it expels the water it goes in the opposite direction.

∴ total momentum = MV - mv = 0

M = mass of the squid without water = 7.50-1.55 = 5.95 kg

V = speed of the squid = 2.6 m/s

m = mass of the water expelled = 1.55 kg

v = speed of the water

=> 5.95 x 2.6 = v x 1.55

=> v = 9.98 m/s

(b) Kinetic energy is the total kinetic energy of both the water and the squid

=> K.E = MV2/2 + mv2/2 = 97.3 J